I Believe Its 17 But Im Not Shure. Please Help?
A rectangle has a length of 15 feet and a width of 8 feet. How long is a diagonal of this rectangle?

Found 2 solutions by mananth, Theo:
Answer by mananth(16949)

diagonal is the hypotenuse and length & width are the legs
15^2+8^2= diagonal ^2
225+64=289= diagonal^2
Diagonal = sqrt 289 = 17 feet

Answer by Theo(13342)

the diagonal forms the hypotenuse of a right triangle.

the length is one leg of the right triangle.

the width is the other leg.

if you let:

a = 1 leg of the right triangle
b = the other leg of the right triangle
c = the hypotenuse of the right triangle

then:

c^2 = a^2 + b^2 which becomes:

c^2 = 15^2 + 8^2 which becomes:

c^2 = 225 + 64 which becomes:

c^2 = 289.

Take the square root of 289 to get:

c = 17.

You got it right.